# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

### v2

@@ Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this problem by estimating the Average Treatment Effect (ATE) from decentralized observational data via a Federated Learning (FL) approach, allowing inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores by computing a federated weighted average of local scores with Membership Weights (MW)—probabilities of site membership conditional on covariates—which can be flexibly estimated using parametric or non-parametric classification models. Unlike density ratio weights (DW) from the transportability and generalization literature, which either rely on strong modeling assumptions or cannot be implemented in FL, MW can be estimated using standard FL algorithms and are more robust, as they support flexible, non-parametric models—making them the preferred choice in multi-site settings with strict data-sharing constraints. The resulting propensity scores are used to construct Federated Inverse Propensity Weighting (Fed-IPW) and Augmented IPW (Fed-AIPW) estimators. Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions. Both theoretical analysis and experiments on simulated and real-world data highlight their advantages over meta-analysis and related methods.
@@ 1 Introduction
The Average Treatment Effect (ATE) is a key causal estimand used to quantify the effect of a treatment on an outcome and is commonly employed as the primary measure of efficacy in evaluating new therapies, including vaccines, before regulatory approval (Polack et al., 2020). In Randomized Clinical Trials (RCTs), treatment assignment is randomized, ensuring that the observed association between treatment and outcome reflects a causal effect. Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023; EMA, 2024; Lei and Ding, 2021). However, RCTs are often expensive, time-consuming, or infeasible. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán, 2018; Hernán and Robins, 2006). Although such real-world data is abundant, drawing causal inferences from it is challenging due to confounding covariates, rendering the unadjusted DM estimator biased (Grimes and Schulz, 2002). Adjusting for confounders is thus essential (VanderWeele, 2019). This can be done by predicting counterfactual outcomes before averaging the differences (the G-formula plug-in estimator, Robins, 1986). Another approach is to weight individuals according to their treatment probability, emulating a randomized trial. For instance, the Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin, 1983) relies on estimating the propensity score—the probability of treatment given covariates. Doubly robust estimators such as the Augmented IPW (AIPW) (Bang and Robins, 2005) combine weighting with outcome modeling to remain consistent as long as either model is correctly specified.
@@
Related work. Federated causal inference is still nascent. Khellaf et al. (2025) estimate federated G-formula ATE estimates across multiple RCTs by fitting parametric outcome models at each site via FL, achieving lower variance than DM. In observational settings, Vo et al. (2022) propose a federated Bayesian approach using Gaussian processes with a shared covariance kernel, but it requires sharing the first four moments of the data, limiting scalability and efficiency. Guo et al. (2025) learn a global propensity score via consensus voting over parametric parameters, retaining only sites meeting a shared specification. In contrast, our method assumes no common propensity score: each site may fit its own model.
To address site heterogeneity, Xiong et al. (2023) use a logistic propensity model with shared and site-specific parameters, federating only the common ones. Yin et al. (2025) fit a global model adjusting for covariates and site membership but limit heterogeneity to a site-specific scalar. By contrast, our method makes no structural assumptions, enabling fully nonparametric estimation with heterogeneous local models and relaxing the need for local overlap at each site.
A related body of work focuses on generalizing causal findings from multi-site source populations to a target population. Han et al. (2025) use density ratio weighting of local ATEs to adjust for covariate shift but assume homogeneous nuisance functions across sites and rely on meta-analysis of aggregate statistics. Guo et al. (2024) extend this idea by applying density ratio weights to aggregate local propensity scores to construct a target-specific score, which requires density estimation within each treatment arm at each site—demanding large sample sizes per arm for stable estimates. In both cases, non-parametric density ratio estimation is infeasible under FL constraints, as it requires sharing raw data or detailed covariate representations (e.g., kernel evaluations or high-dimensional histograms). In contrast, our MW-based approach leverages flexible parametric or non-parametric supervised models (e.g., logistic regression, neural networks, gradient-boosted trees) for which federated training is already established, operating without sharing raw data, potentially at lower sample complexity, with full support for heterogeneous nuisance functions, and allowing the natural inclusion of external control arms.
@@ 2.1 ATE Estimators from Centralized Multi-Site Observational Data
$(Y(0),Y(1))\perp\!\!\!\perp H\mid X$.
Robertson et al. (2021) refer to ˜2 as the no center-outcome association condition. It can be stated as a testable null hypothesis requiring that, for every treatment level $w$ and any pair of centers $(k,k^{\prime})$, $\mathbb{E}[Y\mid X,W=w,H=k]=\mathbb{E}[Y\mid X,W=w,H=k^{\prime}]$. Combined with Assumption 1, this ensures that $X$ forms a sufficient set of covariates for confounding adjustment. Our setting is depicted in the graphical model in Figure 1, highlighting that we remove any direct effect of the site on the outcome.
@@
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),\quad\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{*}}=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (1)
where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$. These oracle estimators are unbiased and asymptotically normal.
@@
The above asymptotic variances align with those in the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the (non-oracle) AIPW estimator is inherently doubly robust, remaining consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al., 2018).
@@ 2.2 Meta-Analysis Estimators
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (2)
@@
$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],\quad\quad V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
@@
$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right],\end{array}\right$.
@@ 3.1 Oracle Federated Estimators
As discussed before, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al., 2025), site differences limited to intercept shifts (Yin et al., 2025), or predefined shared structures (Xiong et al., 2023). In practice, treatment assignment frequently varies across sites due to differences in norms, resources, or clinical practices. To properly account for this heterogeneity, the global propensity score must be expressed as a weighted combination of the site-specific scores (see Appendix A.4). A first choice of weights are the density ratio weights (DW) $\omega_{k}^{\mathrm{DW}}$:
$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}^{\mathrm{DW}}(x)}e_{k}(x),$ (3)
where $f_{k}$ and $f$ are the covariate densities locally and globally. Similar weights are used in transportability methods that reweight data to match a target population (Han et al., 2023, 2025; Guo et al., 2024), though here the objective is to recover the global propensity score across the super-population defined by the $K$ participating sites. Unfortunately, DW estimation in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
Instead, we propose the Membership Weights (MW) $\omega_{k}^{\mathrm{MW}}$:
$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle=\omega_{k}^{\mathrm{MW}}(x)}e_{k}(x),$ (4)
@@
Equations 3 and 4 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$. Building on this decomposition, we define our oracle Federated IPW and AIPW estimators (Fed-(A)IPW). We define the Oracle federated estimators as follows:
@@
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],\quad\quad\quad\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
@@
Theorem 5 (Overlap improvement).
@@ 3.2 Federated estimation
Local propensity scores. Each $e_{k}$ can be estimated using any probabilistic binary classifier, either parametric (e.g., logistic regression or neural networks) or non-parametric ( (e.g., generalized random forests, Lee et al., 2010). A key advantage of our approach is flexibility: sites can use different estimation methods tailored to local data or computational constraints. It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
Federated weights: density ratio vs. membership. Parametric density ratio weights $\omega_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{f_{k}(x)}{f(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights. A common choice is to assume parametric covariate distributions (say, Gaussian), estimate $(\hat{\mu}_{k},\hat{\Sigma}_{k})$ locally, and transmit them once to the server. This requires to communicate $O(Kd^{2})$ parameters and is highly sensitive to model misspecification—an issue that becomes critical in high dimensions. Nonparametric density estimation would relax these assumptions but is statistically inefficient and does not yet have practical federated implementations.
In contrast, our membership weights $\omega_{k}^{\mathrm{MW}}(x)=\mathbb{P}(H=k\mid X=x)$ can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities. Such models are readily supported by modern FL algorithms and software libraries. Using the standard FedAvg algorithm (McMahan et al., 2017) requires exchanging $TKP$ floats (training rounds $\times$ sites $\times$ model parameters), which is feasible for models of practical size and a large number of sites. Non-parametric classifiers like random forests (Hauschild et al., 2022) and gradient-boosted trees (Li et al., 2020) have also been adapted to the federated setting.
@@
Our approach can naturally accommodate missing data. Local propensity scores can be estimated consistently with logistic regression or random forests (Jiang et al., 2020; Josse et al., 2024). For membership weights, constant imputation combined with federated random forests provides a consistent solution (Le Morvan et al., 2021), whereas density ratio weights would require adapting more complex federated EM algorithms (Dieuleveut et al., 2021; Marfoq et al., 2021). Finally, doubly robust estimators with missing data can be obtained via non-parametric federated outcome models (Mayer et al., 2020).

### v3

@@ Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this problem by estimating the Average Treatment Effect (ATE) from decentralized observational data via a Federated Learning (FL) approach, allowing inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores via a federated weighted average of local scores using Membership Weights (MW), defined as probabilities of site membership conditional on covariates. MW can be flexibly estimated with parametric or non-parametric classification models using standard FL algorithms. The resulting propensity scores are used to construct Federated Inverse Propensity Weighting (Fed-IPW) and Augmented IPW (Fed-AIPW) estimators. In contrast to meta-analysis methods, which fail when any site violates positivity, our approach exploits heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions. Theoretical analysis and experiments on simulated and real-world data demonstrate clear advantages over meta-analysis and related approaches.
@@ 1 Introduction
The Average Treatment Effect (ATE) is a key causal estimand used to quantify the effect of a treatment on an outcome and is commonly employed as the primary measure of efficacy in evaluating new therapies, including vaccines, before regulatory approval (Polack et al., 2020). In Randomized Clinical Trials (RCTs), treatment assignment is randomized, ensuring that the observed association between treatment and outcome reflects a causal effect. Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023a; EMA, 2024; Lei and Ding, 2021). However, RCTs are often expensive, time-consuming, or infeasible. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán, 2018; Hernán and Robins, 2006). Although such real-world data is abundant, drawing causal inferences from it is challenging due to confounding covariates, rendering the unadjusted DM estimator biased (Grimes and Schulz, 2002). Adjusting for confounders is thus essential (VanderWeele, 2019). This can be done by predicting counterfactual outcomes before averaging the differences (the G-formula plug-in estimator, Robins, 1986). Another approach is to weight individuals according to their treatment probability, emulating a randomized trial. For instance, the Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin, 1983) relies on estimating the propensity score—the probability of treatment given covariates. Doubly robust estimators such as the Augmented IPW (AIPW) (Bang and Robins, 2005) combine weighting with outcome modeling to remain consistent as long as either model is correctly specified.
@@
Related work. Federated causal inference is still nascent. Khellaf et al. (2025) estimate federated G-formula ATE estimates across multiple RCTs by fitting parametric outcome models at each site via FL, achieving lower variance than DM. In observational settings, Vo et al. (2022) propose a federated Bayesian approach using Gaussian processes with a shared covariance kernel, but it requires sharing the first four moments of the data, limiting scalability and efficiency. Guo et al. (2025) learn a global propensity score via consensus voting over parametric parameters, retaining only sites meeting a shared specification. In contrast, our method assumes no common propensity score: each site may fit its own model. To address site heterogeneity, Xiong et al. (2023) use a logistic propensity model with shared and site-specific parameters, federating only the common ones. Yin et al. (2025) fit a global model adjusting for covariates and site membership but limit heterogeneity to a site-specific scalar. By contrast, our method makes no structural assumptions, enabling fully nonparametric estimation with heterogeneous local models and relaxing the need for local overlap at each site.
A related line of work focuses on generalizing causal findings from multi-site source populations to a target population. Han et al. (2025) use density ratio weighting of local ATEs to adjust for covariate shift but assume homogeneous nuisance functions across sites and rely on meta-analysis of aggregate statistics. Guo et al. (2024) extend this approach by applying density-ratio weights to combine local propensity scores into a target-specific score. While their framework can target the super-population of the $K$ sites as a special case, it is significantly more constrained than our approach: it requires access to individual-level target covariates via an external public dataset, reweights each treatment arm separately within each site (needing large per-arm sample sizes), and relies on direct estimation of density ratios either through strongly parametric logistic models—which are difficult to validate and sensitive to misspecification—or via nonparametric approaches (e.g., kernel-based methods) that involve sharing raw covariates, making them incompatible with federated learning.
To summarize, our approach stands out for its flexibility and ease of implementation in a federated context: (i) It allows local propensity scores to take arbitrary forms and vary freely across sites, unlike Xiong et al. (2023); Han et al. (2025), which rely on strong parametric assumptions, sometimes with shared parameters; (ii) It truly satisfies the federated learning requirement that no raw data be shared, unlike density ratio-based methods that are either highly parametric or lack federated algorithms (Yin et al., 2025; Han et al., 2025; Guo et al., 2024); and (iii) It does not require local overlap, unlike Xiong et al. (2023); Han et al. (2025), and to the best of our knowledge, is the first to provably improve overlap relative to meta-analysis.
@@ 2.1 ATE Estimators from Centralized Multi-Site Observational Data
$\{Y(0),Y(1)\}\perp\!\!\!\perp H\mid X.$
We provide a federated test for ˜2 in Section˜F.1. With ˜1, they imply $\{Y(0),Y(1)\}\perp H\mid(X,W)$ (Dawid, 1979, Lemma 4.3). This entails the no center-outcome association condition (Robertson et al., 2021), yielding the testable null hypothesis $\mathbb{E}[Y\mid X,W,H]=\mathbb{E}[Y\mid X,W]$. This ensures that $X$ forms a sufficient set of covariates for confounding adjustment. Our setting is depicted in the graphical model in Figure 1, highlighting that we exclude any direct effect of the site on the outcome. In Section˜3.3, we introduce a relaxation of ˜2 and extend our approach to handle the presence of site effects.
@@
$\hat{\tau}_{\mathrm{IPW}}^{*}$ $=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{IPW}}(X_{i};e),$ (1)
$\hat{\tau}_{\mathrm{AIPW}}^{*}$ $=\frac{1}{n}\sum_{i=1}^{n}\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0}),$ (2)
where $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}\!\left(Y_{i}-\mu_{1}(X_{i})\right)}{e(X_{i})}-\frac{(1-W_{i})\!\left(Y_{i}-\mu_{0}(X_{i})\right)}{1-e(X_{i})}$ and $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$. These estimators are unbiased and asymptotically normal.
@@
The above asymptotic variances align with the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the (non-oracle) AIPW estimator is inherently doubly robust, remaining consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al., 2018).
@@ 2.2 Meta-Analysis Estimators
$\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{IPW}}^{(k)},\quad\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\frac{n_{k}}{n}\hat{\tau}_{\mathrm{AIPW}}^{(k)},$ (3)
@@
$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}$ $=\textstyle\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
$V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}$ $=\textstyle\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
@@
$\left\{\begin{array}[]{l}\!V_{\mathrm{IPW}}^{(k)}\!=\!\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\!\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\!\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}\!=\!\mathbb{V}\big[\tau(X)\!\mid\!H=k\big]\!+\!\mathbb{E}\Big[\left(\frac{(Y\!-\!\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H\!=\!k\Big]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1\!-\!e_{k}(X)}\right)\!\mid\!H\!=\!k\right],\end{array}\right$.
@@ 3.1 Oracle Federated Estimators
As discussed before, existing federated causal inference methods often rely on restrictive assumptions—such as a common propensity score across sites (Guo et al., 2025), site differences limited to intercept shifts (Yin et al., 2025), or predefined shared structures (Xiong et al., 2023). In practice, treatment assignment frequently varies across sites due to differences in norms, resources, or clinical practices. This heterogeneity implies that the global propensity score takes the form of a weighted combination of the site-specific scores, that is $e(X)=\sum_{k=1}^{K}\omega_{k}(X)e_{k}(X)$, see Appendix A.4. These weights $\{\omega_{k}\}_{k\in[K]}$ can be written as density ratio weights (DW):
$e(x)=\sum_{k=1}^{K}\underbrace{\rho_{k}\frac{f_{k}(x)}{f(x)}}_{=\omega_{k}}(x)e_{k}(x),$ (4)
where $f_{k}$ and $f$ are the covariate densities locally and globally. Similar weights are used in transportability methods that reweight data to match a target population (Han et al., 2023, 2025; Guo et al., 2024), though here the objective is to recover the global propensity score across the super-population defined by the $K$ participating sites. Unfortunately, estimating DW $\omega_{k}$ in a federated setting requires modeling the $f_{k}$’s, which entails strong distributional assumptions and becomes challenging in high dimensions.
Instead, we propose to rewrite the $\omega_{k}$’s as Membership Weights (MW) :
$e(x)=\sum_{k=1}^{K}\underbrace{\mathbb{P}(H=k\mid X=x)}_{\displaystyle={\omega_{k}}(x)}e_{k}(x),$ (5)
@@
Equations 4 and 5 enable combining locally estimated propensity scores $e_{k}$ into a global propensity score using globally learned weights $\omega_{k}(x)$. Building on this decomposition, we define our oracle Federated IPW and AIPW estimators (Fed-(A)IPW). We define the Oracle federated estimators as follows:
@@
$\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}],$
$\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{*}]$ $=\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{fed^{*}}}]\leq\mathbb{V}[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}],$
@@
Theorem 5.
We compare the improvement in the overlap
@@ 3.2 Federated Estimation
Local propensity scores. Each $e_{k}$ can be estimated using any probabilistic binary classifier, either parametric (e.g., logistic regression or neural networks) or non-parametric ( (e.g., generalized random forests, Lee et al., 2010). A key advantage of our approach is flexibility: sites can use different estimation methods tailored to local data or computational constraints. It also does not require Assumption 4: local scores may approach 0 or 1 provided the global score remains bounded away from these extremes.This, in particular, enables the integration of external control arms (FDA, 2023b; EMA, 2023), where some sites have $e_{k}(X)=0$ for all control patients yet still contribute to the global analysis.
Federated weights: density ratio vs. membership. Parametric density ratio weights $\hat{\omega}_{k}^{\mathrm{DW}}(x)=\rho_{k}\frac{\hat{f}_{k}(x)}{\hat{f}(x)}$ can be implemented in a one-shot fashion by sharing local density parameters with the server, which then reconstructs the global mixture and computes the weights. A common choice is to assume parametric covariate distributions (say, Gaussian), estimate $(\hat{\mu}_{k},\hat{\Sigma}_{k})$ locally, and transmit them once to the server. This requires to communicate $O(Kd^{2})$ parameters and is highly sensitive to model misspecification—an issue that becomes critical in high dimensions. Nonparametric density estimation would relax these assumptions but is statistically inefficient and does not yet have practical federated implementations.
In contrast, $\hat{\omega}_{k}^{\mathrm{MW}}(x)=\widehat{\mathbb{P}}(H\!=\!k\!\mid\!X\!=\!x)$, our membership weights, can be learned with any probabilistic multiclass classifier trained federatively—for example, logistic regression for simplicity and interpretability, or neural networks to capture complex nonlinearities. Such models are readily supported by modern FL algorithms and software libraries. Using the standard FedAvg algorithm (McMahan et al., 2017) requires exchanging $TKP$ floats (training rounds $\times$ sites $\times$ model parameters), which is feasible for models of practical size and a large number of sites. Non-parametric classifiers like random forests (Hauschild et al., 2022) and gradient-boosted trees (Li et al., 2020) have also been adapted to the federated setting.
@@
Our approach can naturally handle missing data. Local propensity scores can be estimated consistently with logistic regression or random forests (Jiang et al., 2020; Josse et al., 2024). For membership weights, constant imputation combined with federated random forests provides a consistent solution (Le Morvan et al., 2021), whereas density ratio weights would require adapting more complex federated EM algorithms (Dieuleveut et al., 2021; Marfoq et al., 2021). Finally, doubly robust estimators with missing data can be obtained via non-parametric federated outcome models (Mayer et al., 2020).
3.3 Accommodating Site Effects
In practice, the site membership $H$ may affect potential outcomes through systematic differences (e.g., infrastructure, staff expertise, or measurement practices). When populations also differ across sites, ignoring such center effects can induce non-negligible confounding bias, since $H$ becomes associated with both $X$ and $Y$. In this section, we extend our framework to accommodate site effects when site-level covariates $Z$ (e.g., center-specific resources, equipment quality etc.) are observed and adequately capture these differences. We thus propose a relaxation of ˜2.
Assumption 5 (Explained Site Effects).
$\{Y(1),Y(0)\}\perp H\mid(X,Z).$
Under ˜5, the effect of $H$ on $Y$ is fully mediated by $Z$, so that conditional on $(X,Z)$, there is no residual prognostic information carried by $H$. This assumption integrates naturally into our framework by incorporating $Z$ into the membership weights decomposition of Eq. (5):
$e(X,Z)=\sum_{k=1}^{K}\mathbb{P}(H=k\mid X,Z)e_{k}(X).$
In practice, we learn the resulting membership weights $\hat{\omega}_{k}^{\mathrm{MW}}(x,z)\!=\!\widehat{\mathbb{P}}(H\!=\!k\!\mid\!X\!=\!x,Z\!=\!z)$ as described in Section 3.2, with $Z$ included as additional covariates. Local propensity scores $e_{k}(X)$ remain functions of $X$ only because $Z$ is constant within each site. The AIPW outcome regressions are likewise extended to $\mu_{w}(X,Z)$.